# NashOpt - A Python Library for Computing Generalized Nash Equilibria

### 07 Mar 2026 v2

@@ 4.2.1 Proximal ADMM solver
For variational LQ games with fixed parameter $p=p_{0}$ and no game objective, we also consider the proximal-ADMM algorithm in [7] as an alternative approach to compute a vGNE. To apply the method, that only handles equality constraints, we reformulate the problem as the following LQ game with $N+1$ players:
@@
s.t. $\bar{A}x+s=\bar{b},\quad A_{\mathrm{eq}}x=b_{\mathrm{eq}}$
$x_{-i}=x_{-i}^{\star},s=s^{\star},\quad i=1,\ldots,N$
@@ 5.1 Game-theoretic linear quadratic regulation
The equilibrium can be computed as the solution of a Nash equilibrium problem by using the nonlinear least-squares approach (6) (without parameter $p_{0}$), using the centralized LQR solution $K_{\rm LQR}$ associated with $(A,B,\sum_{i}Q_{i},\mathrm{blkdiag}(R_{1},\dots,R_{N}))$ as initial guess.
Alternatively, the coupled discrete-time algebraic Riccati equations can be solved using Riccati-based iterations as described in [33, Section III.B] until convergence is reached within a given tolerance.
@@ 5.2 Game-theoretic model predictive control
In (20), $Q_{y,i}\succeq 0$, $Q_{\Delta u,i}\succ 0$, and $\epsilon_{i}\geq 0$ is a slack variable used to soften shared output constraints (with linear penalty $q_{\epsilon,i}\geq 0$). For example, if an agent is interested in controlling only a subset $I_{i}$ of the components of the output vector with unit weights, we have $Q_{y,i}=\sum_{j\in I_{i}}e_{j}e_{j}^{\top}$.
@@ 6 Examples
We report several numerical tests to illustrate the methods illustrated in the previous sections for computing and designing generalized Nash equilibria. All tests are run on a Macbook Pro with Apple M4 Max chip and 64 GB RAM using JAX [8] for automatic differentiation and just-in-time compilation of the involved functions.
@@ 6.1 Linear-quadratic game
These correspond to the optimal active constraint combinations $(1)$, $(1,4)$, and $(1,2,4)$, respectively. By imposing the additional constraint that the Lagrange multipliers associated with the shared constraints are the same for all agents, the vGNE
[…]
is found, corresponding to the active constraint combination $(1,4)$. The CPU time for computing each equilibrium ranges between $3.6$ and $10.3$ ms for the non-variational case and is $17.7$ ms for the variational one using the HiGHS solver [26].
For comparison, we also use the nonlinear least-squares approach (6) to compute the vGNE
[…]
in $233.9$ ms (11 LM iterations) starting from the initial guess $x^{0}=0$. Note that infinitely many different GNEs can exist for the same combination of active constraints, as clearly shown in [24]. The same vGNE is found by running the proximal ADMM approach described in Section 4.2.1, where we set $\gamma=1$, $\rho=1$, $x^{0}=0$, $\lambda^{0}=0$, which takes $14.6$ ms (229 ADMM iterations).
Next, we further compare the MILP, LM, and proximal ADMM approaches on random GNE problems of increasing size. We consider linear-quadratic games with $N$ agents, each controlling $n_{i}=2$ decision variables, with cost functions $f$ as in (21) and shared constraints defined by $n_{g}=2N$ random linear inequalities with unit right-hand side. The results are shown in Figure 1(a), where, for an increasing number $N$ of agents, we report the CPU time for computing a GNE via MILP and for computing a vGNE by LM and proximal ADMM. Computing GNEs on LQ games via MILP with HiGHS is about two orders of magnitude more efficient than by LM, and even more by using Gurobi’s the state-of-the-art MILP solver; it also outperforms proximal ADMM (which takes between 258 and 1794 iterations). Figure 1(b) shows instead the CPU time taken when fixing $N=5$ agents and increasing the number of shared constraints $n_{g}$ up to 300, showing that MILP methods degrade more in performance compared to proximal ADMM due to the increasing number of binary variables in (14).
@@ 6.2 Inverse linear-quadratic game design
We consider an inverse game-design problem for a GNEP with $N=10$ agents, each optimizing a vector $x_{i}\in\mathbb{R}^{n_{i}}$ of $n_{i}=10$ variables ($x\in\mathbb{R}^{100})$, parameterized by a vector $p\in\mathbb{R}^{n_{p}}$, with $n_{p}=5$, with $\|p\|_{\infty}\leq 100$. For a given $p$, each agent $i=1,\ldots,N$ solves the convex quadratic optimization problem
$\begin{array}[]{rl}x_{i}^{\star}(p)=\arg\min_{x_{i}}&\frac{1}{2}x^{\top}Q^{i}x+(c^{i}+F^{i}p)^{\top}x\\ \text{s.t. }&Ax\leq b+Sp\\ &A_{\mathrm{eq}}x=b_{\mathrm{eq}}+S_{\mathrm{eq}}p\\ &\ell_{i}\leq x\leq u_{i}\\ &x_{-i}=x^{\star}_{-i}(p)\\ &i=1,\dots,N\end{array}$ (22)
with $b\in\mathbb{R}^{50}$, $b_{\mathrm{eq}}\in\mathbb{R}^{5}$. All matrices and vectors are randomly generated with appropriate dimensions, with $Q^{i}\succ 0$, and the box constraints are defined by $u=-\ell=10\cdot\mathbf{1}$
To formulate the inverse game problem in an ideal setting, we first compute a reference equilibrium $x_{\rm des}$ and a corresponding parameter vector $\hat{p}$ by solving (11) with zero objective $J(x^{\star},p)=0$. Then, we solve the inverse game-design problem (14) by using either $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{\infty}$ (MILP problem) or $J(x^{\star},p)=\frac{1}{2}\|x^{\star}-x_{\rm des}\|_{2}^{2}$ (MIQP problem) to retrieve a parameter vector $p^{\star}$ that yields an equilibrium $x^{\star}(p^{\star})\approx x_{\rm des}$. We use Gurobi in both cases as solver.
@@ 6.3 Stackelberg game
Consider a problem with $N$ agents (the followers), each controlling a decision variable $x_{i}\geq 0$, $i=1,\ldots,N$, subject to the shared constraint $\sum_{i=1}^{N}x_{i}\leq C$. Each follower $i$ minimizes:
@@ 6.4 LQR game
We consider a linear system as in (18) with $n_{x}=10$ states and $n_{u}=10$ inputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,\ldots,N$, $N=10$ agents), with $A,B\in\mathbb{R}^{10\times 10}$ randomly generated and scaled so that $A$ is unstable with spectral radius equal to 1.1. We consider state-weighting matrices $Q_{i}=e_{i}e_{i}^{T}$ and $R_{i}\equiv 10$.
In (19), we set $N_{\textrm{L}QR}=50$, which is a long-enough horizon length used to approximate the infinite-horizon cost, therefore providing a good approximation of the solution of the algebraic Riccati equation associated with the quadruple $((A-B_{-i}K_{-i}),B_{i},Q_{i},R_{i})$ for each agent $i$.
The problem is solved as in (6) (without parameter $p_{0}$) in 6.05 s (6 LM iterations), leading to an asymptotically closed-loop matrix $A-BK^{\star}$ with spectral radius 0.7525. For comparison, the centralized (cooperative) LQR solution leads to a spectral radius of 0.3965. Solving the same problem by using the Riccati-based iterations described in [33, Section III.B] leads to the same solution in 0.11 s (7 Riccati iterations).
@@ 6.5 Game-theoretic MPC
We consider a linear system as in (18) with $n_{x}=6$ states, $n_{u}=3$ inputs, and $n_{y}=3$ outputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,2,3$, $N=3$ agents), with $A,B\in\mathbb{R}^{6\times 3}$ randomly generated and scaled so that $A$ is stable with spectral radius equal to 0.95. The output matrix $C\in\mathbb{R}^{3\times 6}$ is scaled so that the steady-state gain from inputs to outputs is the identity matrix. Each agent $i$ has state-weighting matrix $Q_{i}=C^{\top}Q_{y,i}C$, with $Q_{y,i}=e_{i}e_{i}^{\top}$, $Q_{y,i}\in\mathbb{R}^{3\times 3}$, and input-increment weighting $Q_{\Delta u,i}=0.5$ for $i=1,2,3$. The prediction horizon is $T=10$ and the constraint horizon is $T_{c}=3$. Input constraints are $0\leq u_{i}\leq 4$ for all $i=1,2,3$, and output constraints are $0\leq y_{i}\leq 5$ for all $i=1,2,3$. The slack penalty is $q_{\epsilon,i}=10^{3}$ for all $i=1,2,3$.
@@
The CPU time for solving each GNEP ranges between $4.76$ and $87.76$ ms (HiGHS MILP), while the CPU time for solving each centralized MPC problem ranges between $0.42$ and $1.50$ ms (QP).
@@
6.6 Sparse GNE computations
Consider a generalized Nash equilibrium problem (GNEP) with $N=40$ agents, each one controlling a scalar decision variable $x_{i}$. Agents are grouped into pairs (2k1,2k)(2k-1,2k), $k=1,\ldots,20$, with the agents’ objectives defined as
@@ 6.6 Sparse GNE computations
$J(x,p)=\sum_{i=1}^{N}(x_{i}-x_{i}^{\mathrm{ref}})^{2}+\alpha_{1}\|x\|_{1}$
where the components of the reference vector $x^{\mathrm{ref}}_{i}=\frac{\lceil(i+1)/2\rceil}{10}$, $i=1,\ldots,N$, and $\alpha_{1}>0$ promotes sparsity in the equilibrium solution. Note that in this case there are no game parameters and $p$ is omitted.
Figure 4 shows the number of nonzeros in the computed GNE solution $x^{\star}$ and the optimal cost $J(x^{\star})$ as a function of the $\ell_{1}$-regularization parameter $\alpha_{1}$, obtained by solving Problem (10) with $\rho=10^{4}$. The CPU time is 1.4575 s on the first run due to JAX compilation time, and 0.0314 s on average on the subsequent runs. As expected, as $\alpha_{1}$ increases, the number of zeros in $x^{\star}$ increases, leading to sparser Nash equilibria, at the expense of a larger optimal cost $J(x^{\star})$. Clearly, due to the equilibrium condition $x_{2k-1}^{\star}=x_{2k}^{\star}$, the number of nonzeros in $x^{\star}$ can only decrease by two at a time as $\alpha_{1}$ increases.
Figure 4: Number of nonzeros in $x^{*}$ and optimal cost J(x)J(x*) as a function of the $\ell_{1}$-regularization parameter $\alpha_{1}$.
Figure 5 shows the CPU time required to compute the sparse GNE solution as a function of the number $N$ of agents for $\alpha_{1}=2$ using L-BFGS-B. The CPU time grows roughly linearly with $N$.

### 18 Mar 2026 v3

@@ 4.2.1 Proximal ADMM solver
For vGNEs of LQ games with fixed parameter $p=p_{0}$ and no game objective, we also consider the proximal-ADMM algorithm in [7] as an alternative approach to compute a vGNE. To apply the method, that only handles equality constraints, we reformulate the problem as the following LQ game with $N+1$ players:
@@
s.t. $\bar{A}x+s^{\star}=\bar{b},\quad A_{\mathrm{eq}}x=b_{\mathrm{eq}}$
$x_{-i}=x_{-i}^{\star},\quad i=1,\ldots,N$
@@ 5.1 Game-theoretic linear quadratic regulation
The equilibrium can be computed as the solution of a Nash equilibrium problem by using the nonlinear least-squares approach (6) (without parameter $p_{0}$), using the centralized LQR solution $K_{\rm LQR}$ associated with $(A,B$, $\sum_{i}Q_{i}$, $\mathrm{blkdiag}(R_{1},\dots,R_{N}))$ as initial guess. Note that, in this case, the optimality conditions (4) reduce to $\nabla_{K_{i}}f_{i}(K_{i},K_{-i})=K_{i}^{N_{\rm LQR}}(K_{-i})-K_{i}=0$ for all $i=1,\ldots,N$, where with a slight abuse of notation, we considered $K_{i}$ as its vectorized version; hence, the nonlinear least-squares problem (6) reduces to minimizing the sum of the squared deviations from the optimal gains, i.e., $\sum_{i=1}^{N}\|K_{i}^{N_{\rm LQR}}(K_{-i})-K_{i}\|_{F}^{2}$.
Alternatively, the coupled discrete-time algebraic Riccati equations can be solved using Riccati-based iterations as described in [35, Section III.B] until convergence is reached within a given tolerance.
@@ 5.2 Game-theoretic model predictive control
In (20), $Q_{\Delta u,i}\succ 0$ and ${Q_{y,i}}\succeq 0$; for example, if an agent is interested in controlling only a subset $I_{i}$ of the components of the output vector with unit weights, we have $Q_{y,i}=\sum_{j\in I_{i}}e_{j}e_{j}^{\top}$. The slack variables $\epsilon_{i}\geq 0$ are used to soften the shared output constraints (with linear penalty $q_{\epsilon,i}\geq 0$).
@@ 6 Examples
We report several numerical tests to illustrate the methods described in the previous sections for computing and designing generalized Nash equilibria. All tests are run on a Macbook Pro with Apple M4 Max chip and 64 GB RAM using JAX [8] for automatic differentiation and just-in-time compilation of the involved functions when required.
@@ 6.1 Linear-quadratic game
These correspond to the optimal active constraint combinations $(1)$, $(1,4)$, and $(1,2,4)$, respectively. By imposing the additional constraint that the Lagrange multipliers associated with the shared constraints are the same for all agents, we get the vGNE
$x^{\star}_{v}=[0.3553\ 0.037\ 0.0431\ -1.5324\ -1.4232\ -1.408]^{\top}$}
corresponding to the active constraint combination $(1,4)$. The CPU time for computing each equilibrium ranges between $4.7$ and $10.1$ ms for the non-variational case and is $8.3$ ms for the variational one using the HiGHS solver [27]. Note that infinitely many different GNEs may exist for the same combination of active constraints, as clearly shown in [25].
For comparison, we also use the nonlinear least-squares approach (6) and compute the same vGNE $x^{\star}_{v}$ as above in $229.9$ ms (10 LM iterations) starting from the initial guess $x^{0}=0$. The same vGNE is found again by running the proximal ADMM approach described in Section 4.2.1, where we set $\gamma=1$, $\rho=1$, $x^{0}=0$, $\lambda^{0}=0$, which takes $13.9$ ms (229 ADMM iterations).
Next, we further compare the MILP, LM, and proximal ADMM approaches on random GNE problems of increasing size. We consider linear-quadratic games with $N$ agents, each controlling $n_{i}=2$ decision variables, with cost functions $f$ as in (21) and shared constraints defined by $n_{g}=2N$ random linear inequalities with unit right-hand side. The results are shown in Figure 1(a), where, for an increasing number $N$ of agents, we report the CPU time for computing a GNE via MILP and for computing a vGNE by LM and proximal ADMM. Computing GNEs on LQ games via MILP with HiGHS is about two orders of magnitude more efficient than by LM, and even more by using Gurobi’s state-of-the-art MILP solver; it also outperforms proximal ADMM (which takes between 258 and 1794 iterations). Figure 1(b) shows instead the CPU time taken when fixing $N=5$ agents and increasing the number of shared constraints $n_{g}$ up to 300, showing that MILP methods degrade more in performance compared to proximal ADMM due to the increasing number of binary variables in (14).
@@ 6.2 Inverse linear-quadratic game design
We consider an inverse game-design problem for a GNEP with $N=10$ agents, each optimizing a vector $x_{i}\in\mathbb{R}^{n_{i}}$ of $n_{i}=10$ variables ($x\in\mathbb{R}^{100})$, parameterized by a vector $p\in\mathbb{R}^{n_{p}}$, with $n_{p}=5$, with $\|p\|_{\infty}\leq 100$. For each given $p$, each agent $i=1,\ldots,N$ solves the convex quadratic optimization problem defined in (11) with $b\in\mathbb{R}^{50}$, $b_{\mathrm{eq}}\in\mathbb{R}^{5}$. All matrices and vectors are randomly generated with appropriate dimensions, with $Q^{i}\succ 0$, and the box constraints are defined by $u=-\ell=10\cdot\mathbf{1}$.
To formulate the inverse game problem in an ideal setting, we first compute a reference equilibrium $x_{\rm des}$ and a corresponding parameter vector $\hat{p}$ by solving (11) with zero objective $J(x^{\star},p)=0$. Then, we solve the inverse game-design problem (14) by using either $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{\infty}$ (MILP problem) or $J(x^{\star},p)=\frac{1}{2}\|x^{\star}-x_{\rm des}\|_{2}^{2}$ (MIQP problem) to retrieve a parameter vector $p^{\star}$ that yields an equilibrium $x^{\star}(p^{\star})\approx x_{\rm des}$. We use Gurobi’s solvers in both cases.
@@ 6.3 Stackelberg game
Consider a problem with $N$ agents (the followers), each controlling a decision variable $x_{i}\geq 0$, $i=1,\ldots,N$, subject to the shared constraint $\sum_{i=1}^{N}x_{i}\leq C$. Each agent $i$ minimizes:
@@ 6.4 LQR game
We consider a linear system as in (18) with $n_{x}=10$ states and $n_{u}=10$ inputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,\ldots,N$, $N=10$ agents), with $A,B\in\mathbb{R}^{10\times 10}$ randomly generated and scaled so that $A$ is unstable with spectral radius equal to 1.1. We consider state-weighting matrices $Q_{i}=e_{i}e_{i}^{T}$ and $R_{i}=10$ for all $i=1,\ldots,N$.
In (19), we set $N_{\textrm{L}QR}=50$, which is a long-enough horizon used to approximate the infinite-horizon cost, therefore providing a good approximation of the solution of the algebraic Riccati equation associated with the quadruple $((A-B_{-i}K_{-i}),B_{i},Q_{i},R_{i})$ for each agent $i$.
The problem is solved as in (6) (without parameter $p_{0}$) in 6.05 s (6 LM iterations), leading to an asymptotically closed-loop matrix $A-BK^{\star}$ with spectral radius 0.7525. For comparison, the centralized (cooperative) LQR solution leads to a spectral radius of 0.3965. Solving the same problem by using the Riccati-based iterations described in [35, Section III.B] leads to the same solution in 0.11 s (7 Riccati iterations).
@@ 6.5 Game-theoretic MPC
We consider a linear system as in (18) with $n_{x}=6$ states, $n_{u}=3$ inputs, and $n_{y}=3$ outputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,2,3$, $N=3$ agents), with $A,B\in\mathbb{R}^{6\times 3}$ randomly generated and scaled so that $A$ is stable with spectral radius equal to 0.95. The output matrix $C\in\mathbb{R}^{3\times 6}$ is scaled so that the steady-state gain from inputs to outputs is the identity matrix. Each agent $i$ has state-weighting matrix $Q_{i}=C^{\top}Q_{y,i}C$, with $Q_{y,i}=e_{i}e_{i}^{\top}$, $Q_{y,i}\in\mathbb{R}^{3\times 3}$, and input-increment weighting $Q_{\Delta u,i}=0.5$ for $i=1,2,3$. The prediction horizon is $T=10$ and the constraint horizon is $T_{c}=3$. For all $i=1,2,3$, the input constraints are $0\leq u_{i}\leq 4$ and the output constraints are $0\leq y_{i}\leq 5$, with slack penalty $q_{\epsilon,i}=10^{3}$.
@@
The CPU time for solving each GNEP ranges between $4.36$ and $65.62$ ms (HiGHS MILP), while the CPU time for solving each centralized MPC problem ranges between $0.57$ and $2.80$ ms (QP).
@@
6.6 Computing sparse equilibria
Consider a Nash equilibrium problem with $N=40$ agents, each one controlling a scalar decision variable $x_{i}$ and no free parameter (e.g., $p=0$). The agents are grouped into pairs (2k1,2k)(2k-1,2k), $k=1,\ldots,20$, with the agents’ objectives defined as
@@ 6.6 Computing sparse equilibria
$J(x,0)=\sum_{i=1}^{N}(x_{i}-x_{i}^{\mathrm{ref}})^{2}+\alpha_{1}\|x\|_{1}$
where the components of the reference vector $x^{\mathrm{ref}}_{i}=\frac{\lceil(i+1)/2\rceil}{10}$, $i=1,\ldots,N$, and $\alpha_{1}>0$ promotes sparsity in the equilibrium solution.
Figure 4 shows the number of nonzeros in the computed Nash equilibrium $x^{\star}$ and the optimal cost $J(x^{\star},0)$ as a function of the $\ell_{1}$-regularization parameter $\alpha_{1}$, obtained by solving Problem (10) with $\rho=10^{4}$ and $P=\{0\}$. The CPU time is 1.4350 s on the first run due to JAX compilation time, and 29.9 ms on average on the subsequent runs. As expected, as $\alpha_{1}$ increases, the number of zeros in $x^{\star}$ increases, leading to sparser Nash equilibria, at the expense of a larger optimal cost $J(x^{\star},0)$. Clearly, due to the equilibrium condition $x_{2k-1}^{\star}=x_{2k}^{\star}$, the number of nonzeros in $x^{\star}$ can only decrease by two at a time as $\alpha_{1}$ increases.
Figure 4: Number of nonzeros in $x^{\star}$ and optimal cost $J(x^{\star},0)$ as a function of the $\ell_{1}$-regularization parameter $\alpha_{1}$.
Figure 5 shows the CPU time required to compute the sparse solution as a function of the number $N$ of agents for $\alpha_{1}=2$ using L-BFGS-B. The CPU time grows roughly linearly with $N$.
@@
Acknowledgments
We thank the anonymous reviewers for their constructive comments and suggestions.